Find the probability of getting 2 face cards (king, queen, jack) when 3 cards are drawn from a deck without replacement.
Please help!!

Answer by stanbon(57337)

Find the probability of getting 2 face cards (king, queen, jack) when 3 cards are drawn from a deck without replacement.
-------
There are 12 face cards.
# of ways to succeed: 12C2*40C1 = 66*40 = 2640
----
# of random triples: 52C3 = 22100
----------
Probility = 2640/22100 = 0.1195
==================================
Cheers,
Stan H.